Nightmoon:

The measure of an angle is 5.4°. What is the measure of its supplementary angle?

kacchanforlife:

x = complement x + 5.2 = first angle (5.2 more than complement) measure of two complementary angles add up to 90. x + x+ 5.2 = 90 2x = 84.8 x = 42.4 other angle = x + 5.2 = 42.4 + 5.2 = 47.6 hope this helps! @Dabi

Nightmoon:

Thank you!!

kacchanforlife:

np!